# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ C.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
$|R_{t}$ $(z^{1:D}_{t},z^{1:D}_{t+dt})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})|\leq\sqrt{2\;D\;\sum^{D}_{d=1}C^{2}_{z^{d}}\sum_{z^{d}_{1}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log\frac{p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}{p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}}$
$\leq\sqrt{2\;D\;\underset{d\in\{1,\ldots,D\}}{\sup}\{C^{2}_{z^{d}}\}\sum^{D}_{d=1}\sum_{z^{d}_{1}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\left[\log p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})-\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right]}$
$=\sqrt{C_{2}\left(C_{3}-\underbrace{\sum^{D}_{d=1}\sum_{z^{d}_{1}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}\right)},$
where in the last step we rearrange the terms independent of the approximation parametrized by $\theta$ as constants: $C_{2}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{d}}\}$, $C_{3}=\sum_{z^{d}_{1}\in\mathcal{Z}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$. We now develop the underbraced term in the last equation [^4] In this step, we omit some subscripts from joint probability distributions as they are not defined in the main paper, but they can be inferred from context.:
$\sum^{D}_{d=1}\sum_{z^{d}_{1}}$ $p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})=$
$=\sum^{D}_{d=1}\sum_{z^{d}_{1}}\frac{p(z^{d}_{1},z^{1:D}_{t})}{p_{t}(z^{1:D}_{t})}\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{d}_{1}}\sum_{z_{1}^{1:D\backslash d}}p(z^{d}_{1},z^{1:D}_{t},z^{1:D\backslash d}_{1})\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{1:D}_{1}}p(z^{1:D}_{1},z^{1:D}_{t})\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{1:D}_{1}}p_{1}(z^{1:D}_{1})p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum_{z^{1:D}_{1}}p_{1}(z^{1:D}_{1})p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$
$=-C_{4}\;\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\underbrace{\sum^{D}_{d=1}-\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}_{\textrm{Cross-entropy}}\right],$
where $C_{4}=1/p_{1}(z^{1:D}_{1})$.
@@
$\left|R_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})\right|^{2}\leq C_{2}C_{3}+C_{2}C_{4}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right],$
@@
C.1.2 Bounded Deviation of the Generated Distribution
As proceeded in Sec. C.1.1, we start by introducing the necessary concepts that will reveal useful for the proof of the intended result.
@@ C.1.2 Bounded Deviation of the Generated Distribution
where $R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ denotes the unconditional multivariate rate matrix defined in Eq. 13. This process can be simulated exactly using Gillespie’s Algorithm (Gillespie, 1976; 1977). However, such an algorithm does not scale for large $D$ (Campbell et al., 2022). Although $\tau$-leaping is a widely adopted approximate algorithm to address this limitation (Gillespie, 2001), it requires ordinal discrete state spaces, which is suitable for cases like text or images but not for graphs. Therefore, we cannot apply it in the context of this paper. Additionally, directly replacing the infinitesimal step $\text{d}t$ in Eq. 13 with a finite time step $\Delta t$ la Euler method is inappropriate, as $R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ prevents state transitions in more than one dimension per step under the continuous framework. Instead, Campbell et al. (2024) propose an approximation where the Euler step is applied independently to each dimension, as seen in Eq. 4.
@@
[…]
$\mathcal{R}(t,t+\Delta t)=e^{R\Delta t}$ (19)
@@
For $z_{t}^{1:D}\in\mathcal{Z}^{D},z_{1}^{d}\in\mathcal{Z}$ and $t\in[0,1]$, for each variable $z^{d}$ of a joint variable $z^{1:D}$, there exists a constant $B^{d}_{t}>0$ such that $p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\leq B^{d}_{t}\,p^{d}_{t|1}(z^{d}_{t}|z^{d}_{1})^{2}$.
@@
with $B_{t}=\underset{d\in{1,\ldots,D}}{\sup}B^{d}_{t}$.
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;R^{d}_{t}(z^{1:D}_{t},z^{d}_{t+\mathrm{d}t})$
$=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\mathbb{E}_{p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})}\left[R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})\right]$
$=\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\sum_{z^{d}_{1}\in\mathcal{Z}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1}).$
@@
$\left|\partial_{t}R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})\right|$ $=\left|\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\sum_{z^{d}_{1}\in\mathcal{Z}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\,\partial_{t}R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\sum_{z^{d}_{1}\in\mathcal{Z}}p^{d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\,\left|\partial_{t}R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\backslash d}_{t},z^{1:D\backslash d}_{t+\mathrm{d}t})\;\sum_{z^{d}_{1}\in\mathcal{Z}}B^{d}_{t}\,p^{d}_{t|1}(z^{d}_{t}|z^{d}_{1})^{2}\frac{2}{p^{d}_{t|1}(z^{d}_{t}|z^{d}_{1})^{2}}$
@@
where in the first inequality triangular we apply triangular inequality; in the second inequality, we use Lemma 5 and in 6 to upper bound $|\partial_{t}R^{*d}_{t}(z^{d}_{t},z^{d}_{t+\mathrm{d}t}|z^{d}_{1})|$ and $p(z^{d}_{1}|z^{1:D}_{t})$, respectively. ∎
Now, we finally start the proof of Theorem 3
@@
Let $\{z^{1:D}_{t}\}_{t\in[0,1]}\in\mathcal{Z}^{D}\times[0,1]$ be a CTMC starting with $p(z^{1:D}_{0})=p_{\epsilon}$ and ending with $p(z^{1:D}_{1})=p_{\text{data}}$, whose groundtruth rate matrix is $R_{t}$. Additionally, let $(y^{1:D}_{k})_{k=0,1,\ldots,K}$ be a Euler sampling approximation of that CTMC, with maximum step size $\Delta T=\sup_{k}\Delta t_{k}$ and an approximate rate matrix $R^{\theta}_{t}$. Then, under 6, the following total variation bound holds:
$\|p(y^{1:D}_{K})-p_{\text{data}}\|_{\text{TV}}\leq UZD\;+\;B(ZD)^{2}\Delta t+\;O(\Delta t),$
where $U=\underset{\begin{subarray}{c}t\in[0,1],\;\\ z^{1:D}_{t},\;z^{1:D}_{t+\mathrm{d}t}\in\mathcal{Z}^{D}\end{subarray}}{\sup}\sqrt{C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right]}$ and $B=\underset{\begin{subarray}{c}t\in[0,1],\,z_{1}^{d}\in\mathcal{Z}\;\\ z_{t}^{1:D}\in\mathcal{Z}^{D}\end{subarray}}{\sup}B^{d}_{t}$.

### v3

@@ D.1.1 Bounded Estimation Error of Unconditional Multivariate Rate Matrix
$|R_{t}$ $(z^{1:D}_{t},z^{1:D}_{t+dt})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})|\leq\sqrt{2\;D\;\sum^{D}_{d=1}C^{2}_{z^{(d)}}\sum_{z^{(d)}_{1}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log\frac{p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}{p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}}$
$\leq\sqrt{2\;D\;\underset{d\in\{1,\ldots,D\}}{\sup}\{C^{2}_{z^{(d)}}\}\sum^{D}_{d=1}\sum_{z^{(d)}_{1}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\left[\log p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})-\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right]}$
$=\sqrt{C_{2}\left(C_{3}-\underbrace{\sum^{D}_{d=1}\sum_{z^{(d)}_{1}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}\right)},$
where in the last step we rearrange the terms independent of the approximation parametrized by $\theta$ as constants: $C_{2}=2D\,{\sup}_{{d\in\{1,\ldots,D\}}}\{C^{2}_{z^{(d)}}\}$, $C_{3}=\sum_{z^{(d)}_{1}\in\mathcal{Z}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$.
We now develop the underbraced term in the last equation [^4] In this step, we omit some subscripts from joint probability distributions as they are not defined in the main paper, but they can be inferred from context.:
$\sum^{D}_{d=1}\sum_{z^{(d)}_{1}}$ $p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})=$
$=\sum^{D}_{d=1}\sum_{z^{(d)}_{1}}\frac{p(z^{(d)}_{1},z^{1:D}_{t})}{p_{t}(z^{1:D}_{t})}\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{(d)}_{1}}\sum_{z_{1}^{1:D\setminus(d)}}p(z^{(d)}_{1},z^{1:D}_{t},z^{1:D\setminus(d)}_{1})\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{1:D}_{1}}p(z^{1:D}_{1},z^{1:D}_{t})\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum^{D}_{d=1}\sum_{z^{1:D}_{1}}p_{1}(z^{1:D}_{1})p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$
$=\frac{1}{p_{t}(z^{1:D}_{t})}\sum_{z^{1:D}_{1}}p_{1}(z^{1:D}_{1})p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$
$=-C_{4}\;\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\underbrace{\sum^{D}_{d=1}-\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}_{\textrm{Cross-entropy}}\right],$
where $C_{4}=1/p(z^{1:D}_{1})$.
@@
$\left|R_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})-R^{\theta}_{t}(z^{1:D}_{t},z^{1:D}_{t+dt})\right|^{2}\leq C_{2}C_{3}+C_{2}C_{4}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right],$
@@
Corollary 0 (Bounded estimation error of unconditional rate matrix for graphs).
Given $t\in[0,1]$, graphs $G_{t},G_{t+\mathrm{d}t}$, and $G_{1}\sim p_{1}(G_{1})$, let $R_{t}(G_{t},G_{t+\mathrm{d}t})$ be the groundtruth rate matrix of the CTMC, which we approximate with $R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})$. Then, there exist constants $\bar{C}_{0},\bar{C}_{1},\bar{C}_{3}$ such that:
$|R_{t}(G_{t},G_{t+\mathrm{d}t})-$ $R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})|^{2}\leq$
$\leq$ $\bar{C}_{0}+\bar{C}_{1}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum^{N}_{n=1}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]$
$+\bar{C}_{2}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{1\leq i<j\leq N}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right].$
Proof.
Recall that a graph $G$ is defined as the set of nodes and edges $G=(x^{1:n:N},e^{1:i<j:N})$, such that $x^{(n)}\in\mathcal{X}=\{1,\dots,X\}$, $e^{(ij)}\in\mathcal{E}=\{1,\dots,E\}$, so we can develop Eq. 11 as follows:
$R_{t}(G_{t},G_{t+\mathrm{d}t})=$ $\underbrace{\sum_{n=1}^{N}\delta\left(G^{\setminus(n)}_{t},G^{\setminus(n)}_{t+\mathrm{d}t}\right)\;\mathbb{E}_{p_{1|t}(x^{(n)}_{1}|G_{t})}\left[R^{(n)}_{t}(x^{(n)}_{t},x^{(n)}_{t+\mathrm{d}t}|x_{1}^{(n)})\right]}_{R_{t}(x_{t}^{1:n:N},x_{t+\mathrm{d}t}^{1:n:N})}$
$+\underbrace{\sum_{1\leq i<j\leq X}\delta\left(G^{\setminus(ij)}_{t},G^{\setminus(ij)}_{t+\mathrm{d}t}\right)\;\mathbb{E}_{p_{1|t}(e^{(ij)}_{1}|G_{t})}\left[R^{(ij)}_{t}(e^{(ij)}_{t},e^{(ij)}_{t+\mathrm{d}t}|e_{1}^{(ij)})\right]}_{R_{t}(e_{t}^{1:i<j:N},e_{t+\mathrm{d}t}^{1:i<j:N})},$ (16)
where $G^{\setminus(n)}$ and $G^{\setminus(ij)}$ denote the whole graph except the node $n$ and the edge connecting node $i$ to node $j$, respectively.
Therefore, the result from Theorem 11 implies:
$|R_{t}(G_{t},G_{t+\mathrm{d}t})$ $-R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})|\leq$
$\leq$ $|R_{t}(x_{t}^{1:n:N},x_{t+\mathrm{d}t}^{1:n:N})-R^{\theta}_{t}(x_{t}^{1:n:N},x_{t+\mathrm{d}t}^{1:n:N})|+|R_{t}(e_{t}^{1:i<j:N},e_{t+\mathrm{d}t}^{1:i<j:N})-R^{\theta}_{t}(e_{t}^{1:i<j:N},e_{t+\mathrm{d}t}^{1:i<j:N})|$
$\leq$ $\sqrt{C_{0}^{\mathcal{X}}+C_{1}^{\mathcal{X}}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum^{N}_{n=1}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]}$
$+\sqrt{C_{2}^{\mathcal{E}}+C_{3}^{\mathcal{E}}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{1\leq i<j\leq N}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right]}$
We now apply the identity $\sum_{d=1}^{D}\sqrt{x_{d}}\leq\sqrt{D\sum^{D}_{d=1}x_{d}}$, derived from the Cauchy-Schwarz inequality in the proof of Theorem 11, with $D=2$ to obtain:
$|R_{t}(G_{t},G_{t+\mathrm{d}t})-$ $R^{\theta}_{t}(G_{t},G_{t+\mathrm{d}t})|\leq$
$\leq$ $\Bigg{(}\bar{C}_{0}+\bar{C}_{1}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum^{N}_{n=1}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]$
$+\bar{C}_{2}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{1\leq i<j\leq N}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right]\Bigg{)}^{1/2},$
with $\bar{C}_{0}=2(C_{0}^{\mathcal{X}}+C_{2}^{\mathcal{E}})$, $\bar{C}_{1}=2C_{1}^{\mathcal{X}}$, and $\bar{C}_{2}=2C_{3}^{\mathcal{E}}$, yielding the intended result. ∎
D.1.2 Bounded Deviation of the Generated Distribution
As in Sec. D.1.1, we start by introducing the necessary concepts that will reveal useful for the proof of the intended result.
@@ D.1.2 Bounded Deviation of the Generated Distribution
where $R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ denotes the unconditional multivariate rate matrix defined in Eq. 11. This process can be simulated exactly using Gillespie’s Algorithm (Gillespie, 1976, 1977). However, such an algorithm does not scale for large $D$ (Campbell et al., 2022). Although $\tau$-leaping is a widely adopted approximate algorithm to address this limitation (Gillespie, 2001), it requires ordinal discrete state spaces, which is suitable for cases like text or images but not for graphs. Therefore, we cannot apply it in the context of this paper. Additionally, directly replacing the infinitesimal step $\text{d}t$ in Eq. 11 with a finite time step $\Delta t$ la Euler method is inappropriate, as $R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ prevents state transitions in more than one dimension per step under the continuous framework. Instead, Campbell et al. (2024) propose an approximation where the Euler step is applied independently to each dimension, as seen in Eq. 4.
@@
For this proof, we also introduce the notion of Markov kernel of a CTMC. The Markov kernel, $\mathcal{R}_{t\to t+\Delta t}$, is a function that provides the transition probabilities between states over a finite time interval, $\Delta t$:
$p_{t+\Delta t}=\mathcal{R}_{t\to t+\Delta t}^{\top}\,p_{t}\,.$ (18)
For example, for a univariate CTMC with a state space $\mathcal{Z}$ of cardinality $Z$, the Markov kernel $\mathcal{R}_{t\to t+\Delta t}$ is a matrix where each entry $\mathcal{R}_{t\to t+\Delta t}^{(ij)}$ represents the probability that the single variable transitions from state $i$ at time $t$ to state $j$ at time $t+\Delta t$, i.e., $p_{t+\Delta t|t}{(j|i)}$.
The definition of Markov kernel contrasts with the one of rate matrix (or generator), $R_{t}$, which instead characterizes the infinitesimal transition rates between states at a given time $t$. Thus, while the rows of the rate matrix sum to 0, the Markov kernel matrices are stochastic, i.e., $\sum_{j\in Z}\mathcal{R}_{t\to t+\Delta t}^{(ij)}=1,\;\forall i$. This contrasts with the rate matrix where rows sum to 1. Additionally, Markov kernels must also respect the initial condition $\mathcal{R}_{t\to t}=I$, where $I$ denotes the identity matrix.
Recall from Sec. 2 that the evolution of a CTMC is governed by the rate matrix through the equation:
$\partial_{t}p_{t}=R^{\top}p_{t},$
which represents a first-order differential equation. Here, we focus on the time-homogeneous case, where $R_{t}=R$ in the time interval $[t;\,t+\Delta t]$, i.e., the rate matrix remains constant within the time interval. In that case, its solution is given by:
$p_{t+\Delta t}=e^{R^{\top}\Delta t}\,p_{t}$
Therefore the result above sets, by definition, the corresponding Markov kernel of a constant rate matrix in a finite time interval $\Delta t$ as:
$\mathcal{R}_{t\to t+\Delta t}=({e^{R^{\top}\Delta t}})^{\top}=e^{R\Delta t},$ (19)
where the second equality is a direct consequence of the definition of the matrix exponential as a series expansion.
@@
For $z_{t}^{1:D}\in\mathcal{Z}^{D},z_{1}^{(d)}\in\mathcal{Z}$ and $t\in[0,1]$, for each variable $z^{(d)}$ of a joint variable $z^{1:D}$, there exists a constant $B^{(d)}_{t}>0$ such that $p_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\leq B^{(d)}_{t}\,p_{t|1}(z^{(d)}_{t}|z^{(d)}_{1})^{2}$.
@@
with $B_{t}=\underset{d\in{1,\ldots,D}}{\sup}B^{(d)}_{t}$.
@@
$R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})$ $=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;R^{(d)}_{t}(z^{1:D}_{t},z^{(d)}_{t+\mathrm{d}t})$
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\mathbb{E}_{p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})}\left[R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})\right]$
$=\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\sum_{z^{(d)}_{1}\in\mathcal{Z}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1}).$
@@
$\left|\partial_{t}R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})\right|$ $=\left|\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\sum_{z^{(d)}_{1}\in\mathcal{Z}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\,\partial_{t}R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\sum_{z^{(d)}_{1}\in\mathcal{Z}}p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\,\left|\partial_{t}R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})\right|$
$\leq\sum^{D}_{d=1}\delta(z^{1:D\setminus(d)}_{t},z^{1:D\setminus(d)}_{t+\mathrm{d}t})\;\sum_{z^{(d)}_{1}\in\mathcal{Z}}B^{(d)}_{t}\,p_{t|1}(z^{(d)}_{t}|z^{(d)}_{1})^{2}\frac{2}{p_{t|1}(z^{(d)}_{t}|z^{(d)}_{1})^{2}}$
@@
where in the first inequality triangular we apply triangular inequality; in the second inequality, we use Lemma 5 and 6 to upper bound $|\partial_{t}R^{*(d)}_{t}(z^{(d)}_{t},z^{(d)}_{t+\mathrm{d}t}|z^{(d)}_{1})|$ and $p_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$, respectively. ∎
Now, we finally start the proof of Theorem 8.
@@
Let $\{z^{1:D}_{t}\}_{t\in[0,1]}\in\mathcal{Z}^{D}\times[0,1]$ be a CTMC starting with $p_{0}(z^{1:D}_{0})=p_{\epsilon}$ and ending with $p_{1}(z^{1:D}_{1})=p_{\text{data}}$, whose groundtruth rate matrix is $R_{t}$. Additionally, let $(y^{1:D}_{k})_{k=0,\frac{1}{K},\ldots,1}$ be a Euler sampling approximation of that CTMC, with maximum step size $\Delta t=\sup_{k}\Delta t_{k}$ and an approximate rate matrix $R^{\theta}_{t}$. Then, under 6, the following total variation bound holds:
$\|p(y^{1:D}_{1})-p_{\text{data}}\|_{\text{TV}}\leq UZD\;+\;B(ZD)^{2}\Delta t+\;O(\Delta t),$
where $U=\underset{\begin{subarray}{c}t\in[0,1],\;\\ z^{1:D}_{t},\;z^{1:D}_{t+\mathrm{d}t}\in\mathcal{Z}^{D}\end{subarray}}{\sup}\sqrt{C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right]}$ and $B=\underset{\begin{subarray}{c}t\in[0,1],\,z_{1}^{(d)}\in\mathcal{Z}\;\\ z_{t}^{1:D}\in\mathcal{Z}^{D}\end{subarray}}{\sup}B^{(d)}_{t}$.